### v1

@@ Abstract
The increasing size of the Key-Value (KV) cache during the Large Language Models long-context inference is the main obstacle for its balance between the deployment cost and task accuracy. To reduce the KV cache size in such scenarios, most previous efforts leveraged on the attention weight to evict non-critical cache tokens. But there is a trade-off in those methods, they usually require major modifiation of the inference infrastructure and significant computation overhead. Base on the fact that the Large Lanuage models are autoregresssive models, we propose LagKV, a KV allocation strategy only relying on straight forward comparison among KV themself. It is a totally attention free method which offers easy integration to the main stream inference platform and comparable performance comparing to other complicated KV compression methods. Results on LongBench and PasskeyRetrieval show that, our approach achieves nearly zero loss when the ratio is 2×2$\times$2 × and $\approx 90\%$ of the original model performance for 8×8$\times$8 ×. Especially in the 64-digit passkey retrieval task, our mehod outperforms the attention weight based method $H_{2}O$ over $60\%$ with same compression ratios. Our code is available at https://github.com/AI-Lab-China-Merchants-Bank/LagKV.
@@ 1 Introduction
Large Language Models (LLMs) have recently demonstrated remarkable success across diverse text processing tasks, including document retrieval Laban et al. (2023), code generation Gu (2023), and mathematical reasoning (like R1 model DeepSeek-AI et al. (2025)). The Scaling law Kaplan et al. (2020) suggests that larger models generally achieve superior performance. The R1 model further indicates that longer generation sequences with additional ’thinking tokens’ can enhance reasoning capabilities. However, these improvements comes at a significant cost: the growing KV cache size poses a major challenge for efficient LLM inference. Many efforts try to mitigate this challenge.
@@
Most of LLMs are totally relying on Self-Attention mechanism to determine which historical tokens are important in the next token prediction. Therefore, many KV compression approaches are based on it to drop unimportant ones  (Zhang et al., 2024; Xiao et al., 2023; Liu et al., 2024b; Li et al., 2024; Feng et al., 2024). This kind of algorithms keeps a remarkable performance even when the compression ratio is high. However, most of these importance-based token-dropping approaches are incompatible with Flash Attention(FA) because they require computing attention weights to determine token importance. This limitation makes them impractical for deployment, given FA’s critical role in long-context inference.
@@
Another prominent direction in KV cache optimization involves quantization techniques  (Yang et al., 2024; Liu et al., 2024c), which aim to compress the memory footprint of key-value (KV) states by representing them with reduced precision. These methods achieve significant memory savings—often by 4× or more—while preserving model performance through careful error mitigation strategies. Beyond memory efficiency, quantization also reduces the bandwidth overhead of transferring KV cache across devices in distributed inference scenarios, accelerating multi-GPU or memory-bound workloads. However, a critical limitation of pure quantization approaches is that they retain all historical tokens, leaving the computational cost of attention unchanged. For long-context tasks, this means the quadratic complexity of attention persists despite the reduced memory usage.
@@
The simple but with limited performance methods are usually based on the sliding window tokens eviction. Sliding window-based eviction methods—such as those used in Longformer Beltagy et al. (2020), Infinite-LLM Han et al. (2024), and StreamingLLM Xiao et al. (2023)—retain only the initial cache tokens and those within a fixed sliding window, discarding the rest. However, this indiscriminate eviction strategy often leads to a notable degradation in generation quality.
@@
Recent work by  (Liu et al., 2024a; c) addresses the statistical properties of KV states, revealing distinct distribution patterns for keys and values. Their findings suggest that per-channel quantization for keys (which exhibit consistent variance across feature dimensions) and per-token quantization for values (which vary more significantly across sequence positions) yield better fidelity. This observation motivates our key insight: token importance for eviction—traditionally derived from attention weights—can instead be inferred from token- and channel-wise distribution patterns in the KV space. By leveraging these structural properties, we can design a pruning criterion, LagKV, that is both hardware-friendly (compatible with FA) and semantically aware, enabling compute savings alongside memory reduction.
@@
Figure 1: LagKV recursively compression process: partition the KV cache and use the next join chunk as reference to compress the current one. Keep the rest of them as the sliding winow.
@@ 2.1 Preliminaries
LLMs’ next token prediction relies on the previous tokens. First, in the prefill stage, the model uses its tokenizer to convert the words to $n$ indices of the embedding metrics $E\in\mathbb{R}^{V\times d}$ of the model and collects the representations to form a input matrix, $X\in\mathbb{R}^{n\times d}$. This matrix is the initial tokens of the first layer of LLM and then each layer will output a same shape matrix as next layer’s input. To depict the operations in each layer, we follow the notation system from  Liu et al. (2023) with $h$ attention heads. For each head $i\in[1,h]$ and head dimension $d_{h}$, we focus on the Query, Key, and Value states, which are converted from tokens by three linear transformation matrices $W_{i}^{Q}$, $W_{i}^{K}$, $W_{i}^{V}\in\mathbb{R}^{d\times d_{h}}$ sperately:
@@
$Y=Concat_{i\in[1,h]}(A_{i}V_{i})W^{O}\\ \>\text{where}\>A_{i}=\text{softmax}(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{h}}})$ (2)
@@
$K_{i}=Cat[K_{i}:k_{i}],V_{i}=Cat[V_{i}:v_{i}],\\ A_{i}=\text{softmax}(\frac{q_{i}K_{i}^{T}}{\sqrt{d_{h}}})$ (4)
@@ 2.2 LagKV
Since the intrinsic property of autoregressive model, the next generated token will not change abruptly from the previous one. As observed in Liu et al. (2024a), the called token-wise locality will show that the tokens in closer proximity have more similar K/V tensor values compared to tokens that are further apart. And also, the StreamingLLM method Xiao et al. (2023) has already proved that the head part and the sliding window of KV cache are very important. Inspired by these, we proposed our LagKV method as:
@@
  • Always keep the attention sink with size $S$.
  • Partition the rest KV after the sink part with lag size $L$. If it’s not divisible by $L$, the modulo of it will be added to the sliding window.
  • Recursively compute the KV cache score. Use the next partition as a reference, calculate token-wise max and min from the reference then use max-min to normalize the Key and Value states respectly. After the KV are normalized, calculate the channel-wise standard deviation then softmax. The equations are formally like:
    $min_{i}^{p,K,V}=min_{seq}({K_{i}^{p+1},V_{i}^{p+1}})$ (5)
    $max_{i}^{p,K,V}=max_{seq}({K_{i}^{p+1},V_{i}^{p+1}})$ (6)
    $\bar{K_{i}^{p}},\bar{V_{i}^{p}}=\frac{{K_{i}^{p},V_{i}^{p}}-min_{i}^{p,K,V}}{max_{i}^{p,K,V}-min_{i}^{p,K,V}}$ (7)
    $score(K_{i},V_{i})=Softmax(Std.(\bar{K_{i}},\bar{V_{i}}))$ (8)
    where $p$ denotes the partition index, $i$ represents the head index and $seq$ for the sequence axis. Since the last partition has no reference can be used. It will serve as another part of the sliding window.
@@
  • Base on the $score_{i}$, use the Top-k strategy to select tokens in each partition and each head.
The min-max normalization is applied along the sequence dimension, meaning each channel is normalized using statistics from lag-$L$ tokens. Due to token-wise locality, the channel-specific norms of $K_{i}$ and $V_{i}$ are largely eliminated. The resulting normalized representations, $\bar{K_{i}}$ and $\bar{V_{i}}$, retain the original channel-wise variance, allowing the standard deviation to serve as a measure of token importance. The softmax operation then identifies and separates outliers, while the summed scores $score(K_{i})$ and $score(V_{i})$ determine their relative contributions.
As showed in Fig. 1, our method is recursively compressing KV cache in both prefill and decode stages, which is essential for the token-wise locality as mentioned above. It requires relative short distance to keep the similarity among the KV states. Subsequently, another benefit, it also avoids the bias from the long context with length much larger than $L$ and the case when the question is at the end of the prompt Yuan et al. (2024); Tang et al. (2024).
We do not compare the LagKV score to the attention weights here. The attention weights vary on different incoming queries. But our scoring method does not depend on the query states or the tokens after the next joint chunk. It mainly finds the tokens that are not coherent to the next chunk and keep them in the cache. Just like the quantization in KIVI Liu et al. (2024c), we need a rightful mean to find the correct variance and then prune the small ones. However, we use this strategy to evict tokens instead of quantizing them.
To caculate the compresstion ratio, we set the retained token ratio as $r$ in each partition. In the partition chunk, only $rL$ tokens will be kept and others are evicted. Therefore, the compression ratio $C$ for the token sequence length $L_{s}$ not less than $S+2L$ can be expressed as:
$L_{R}=S+rL(Floor(((L_{s}-S)/L)-1)+L+Mod(L_{s}-S,L)$ (10)
@@
Where $L_{R}$ is the length of the KV cache after compression. For the case $L_{s}\leq S+2L$, the compression ratio is zero.
3 Experiments
3.1 Settings
@@
Base Models. We employ two open-source base models: Llama-3.1-8B-Instruct Grattafiori et al. (2024) and Qwen2.5-7B-Instruct Jiang et al. (2023). These models are main stream LLMs with moderate size and both leverage the GQA Ainslie et al. (2023) technique to reduce the KV cache size.
@@
Datasets. We use the facility in Yuan et al. (2024) to extensive test our method. It mainly contains two benchmarks: LongBench Bai et al. (2023) and Needle-in-a-HaystackTest with Passkey-Retrieval Kamradt (2023); Mohtashami & Jaggi (2023). We only test the 64-digit passkey retrieval task which is much more challenging and use the partial match score in the report.
Parameter Spaces. Across the whole paper, we fix the sink size to $S=16$ and vary the lag size $L$ and ratio $r$. The values of $L$ will be $L=128,512,1024$. The values of $r$ will be 2×2$\times$2 ×, 4×4$\times$4 ×, 6×6$\times$6 × and 8×8$\times$8 × which correspond to $r=0.5,0.25,0.167,0.125$ respectly.
Main results. The main results of this work is Table 1.
@@
Llama-3.1-8B-Instruct Baseline 40.71 37.90 28.29 68.49 68.00 58.70 47.44 99.44
@@
Qwen-2.5-7B-Instruct Baseline 41.62 45.00 26.41 68.91 100.00 63.60 51.53 100.00

### v2

@@ Abstract
The increasing size of the Key-Value (KV) cache during the Large Language Models long-context inference is the main obstacle for its balance between the deployment cost and task accuracy. To reduce the KV cache size in such scenarios, most previous efforts leveraged on the attention weight to evict non-critical cache tokens. But there is a trade-off in those methods, they usually require major modification of the inference infrastructure and significant computation overhead. Based on the fact that the Large Language models are autoregressive models, we propose LagKV, a KV compression strategy only relying on straight forward comparison among KV themselves. It is a totally attention free method which offers easy integration to the main stream inference platform and comparable performance comparing to other complicated KV compression methods. Results on RULER benchmark show that, our approach outperforms SnapKV and StreamingLLM in different compression ratios. Especially in the 64-digit passkey retrieval task, our method outperforms the attention weight based method $H_{2}O$ over $50\%$ with same compression ratios. Our code is available at https://github.com/AI-Lab-China-Merchants-Bank/LagKV.
@@ 1 Introduction
Large Language Models (LLMs) have recently demonstrated remarkable success across diverse text processing tasks, including document retrieval Laban et al. (2023), code generation Gu (2023), and mathematical reasoning (like R1 model DeepSeek-AI et al. (2025)). The Scaling law Kaplan et al. (2020) suggests that larger models generally achieve superior performance. The R1-like models further indicates that longer generation sequences with additional ’thinking tokens’ can enhance reasoning capabilities. However, these improvements comes at a significant cost: the growing KV cache size poses a major challenge for efficient LLM inference. Many efforts try to mitigate this challenge.
@@
Most of LLMs are totally relying on Self-Attention mechanism Vaswani et al. (2023) to determine which historical tokens are important in the next token prediction. Therefore, many KV compression approaches are based on it to drop unimportant ones  (Zhang et al., 2024; Liu et al., 2024b; Li et al., 2024; Tang et al., 2024b; NVIDIA, 2024). This kind of algorithms keeps a remarkable performance even when the compression ratio is high. However, most of these importance-based token-dropping approaches depend on the ending query question (Instruction Dependence) to achieve such a performance  Li et al. (2025); Feng et al. (2024); Tang et al. (2024a).
@@
Another prominent direction in KV cache optimization involves quantization techniques  (Yang et al., 2024; Liu et al., 2024c), which aim to compress the memory footprint of KV states by representing them with reduced precision. These methods achieve significant memory savings—often by 4× or more—while preserving model performance through careful error mitigation strategies. Beyond memory efficiency, quantization also reduces the bandwidth overhead of transferring KV cache across devices in distributed inference scenarios, accelerating multi-GPU or memory-bound workloads. However, a critical limitation of pure quantization approaches is that they retain all historical tokens, leaving the computational cost of attention unchanged. For long-context tasks, this means the quadratic complexity of attention persists despite the reduced memory usage.
@@
The simple but with limited performance methods are usually based on the sliding window tokens eviction. Sliding window-based eviction methods—such as those used in Infinite-LLM Han et al. (2024) and StreamingLLM Xiao et al. (2023)—retain only the initial cache tokens and those within a fixed sliding window, discarding the rest. However, this indiscriminate eviction strategy often leads to a notable degradation in generation quality.
@@
Recent work by  (Liu et al., 2024a, c) addresses the statistical properties of KV states, revealing distinct distribution patterns for keys and values. Their findings suggest that per-channel quantization for keys (which exhibit consistent variance across feature dimensions) and per-token quantization for values (which vary more significantly across sequence positions) yield better fidelity. This observation motivates our key insight: token importance for eviction—traditionally derived from attention weights—can instead be inferred from token- and channel-wise distribution patterns in the KV space. By leveraging these structural properties, we can design a pruning criterion, LagKV, that is both hardware-friendly (compatible with Flash Attention Dao (2023)) and instruction independent, enabling compute savings alongside memory reduction.
@@
Figure 1: LagKV recursively compression process: partition the KV cache and use the next joint chunk as reference to compress the current one. Keep the rest of them as the sliding window.
@@ 2.1 Preliminaries
LLMs’ next token prediction relies on the previous tokens. First, in the prefill stage, the model uses its tokenizer to convert the words to $n$ indices of the embedding metrics $E\in\mathbb{R}^{V\times d}$ of the model and collects the representations to form a input matrix, $X\in\mathbb{R}^{n\times d}$. This matrix is the initial tokens of the first layer of LLM and then each layer will output a same shape matrix as next layer’s input. To depict the operations in each layer, we follow the notation system from  Liu et al. (2023) with $h$ attention heads. For each head $i\in[1,h]$ and head dimension $d_{h}$, we focus on the Query, Key, and Value states, which are converted from tokens by three linear transformation matrices $W_{i}^{Q}$, $W_{i}^{K}$, $W_{i}^{V}\in\mathbb{R}^{d\times d_{h}}$ separately:
@@
$Y=Concat_{i\in[1,h]}(A_{i}V_{i})W^{O}$ (2)
where
$A_{i}=\text{softmax}(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{h}}}).$ (3)
@@
$K_{i}=Cat[K_{i}:k_{i}],V_{i}=Cat[V_{i}:v_{i}]$ (5)
$A_{i}=\text{softmax}(\frac{q_{i}K_{i}^{T}}{\sqrt{d_{h}}})$ (6)
@@ 2.2 LagKV
Since the intrinsic property of autoregressive model, the next token representation will not change abruptly from the previous one. As observed in Liu et al. (2024a), the called token-wise locality will show that the tokens in closer proximity have more similar K/V tensor values compared to tokens that are further apart.
And also, the StreamingLLM method Xiao et al. (2023) has demonstrated that the head portion and sliding window of the KV cache are crucial. This suggests that cache compression should use subsequent tokens to assess whether prior tokens remain in the cache, rather than relying on the competition between them—as done in many attention-weight-based methods.
Inspired by above insights, we proposed our LagKV method as:
@@
  • Always keep the attention sink with size $S$ and the already compressed part if had unchanged.
  • Skip the compression if the length of the rest KV after the static part is less than $2L$, where we denote the lag size as $L$.
  • Partition the rest KV with $L$. If it’s not divisible by $L$, the modulo of it will be added to the sliding window.
  • Recursively compute the KV cache score. Use the next partition as a reference, calculate token-wise max and min from the reference then use max-min to normalize the Key and Value states respectively. After the KV are normalized, calculate the channel-wise standard deviation then softmax. The equations are formally like:
    $min_{i}^{p,Z}=min_{seq}({Z_{i}^{p+1}})$ (7)
    $max_{i}^{p,Z}=max_{seq}({Z_{i}^{p+1}})$ (8)
    $\bar{Z_{i}^{p}}=\frac{{Z_{i}^{p}}-min_{i}^{p,Z}}{max_{i}^{p,Z}-min_{i}^{p,Z}}$ (9)
    $score(Z_{i})=Softmax(Std.(\bar{Z_{i}}))$ (10)
    where $Z$ is one of $\{K,V\}$, $p$ denotes the partition index, $i$ represents the head index and $seq$ for the sequence axis. Since the last partition has no reference can be used, our method will naturally have a sliding window with at least size $L$.
@@
  • Base on the $score_{i}$, use the top-K strategy to select tokens in each partition and each head and add them to the compressed part.
The max-min normalization is applied along the sequence dimension, meaning each channel is normalized using statistics from lag-$L$ tokens. Due to token-wise locality, the channel-specific norms of $K_{i}$ and $V_{i}$ are largely eliminated. The resulting normalized representations, $\bar{K_{i}}$ and $\bar{V_{i}}$, retain the original channel-wise variance, allowing the standard deviation to serve as a measure of token importance. The softmax operation then identifies and separates outliers, while the summed scores $score(K_{i})$ and $score(V_{i})$ determine their relative contributions.
As showed in Fig. 1, our method is recursively compressing KV cache in both prefill and decode parts, which is essential for the token-wise locality as mentioned above. It requires relative short distance to keep the similarity among the KV states. Subsequently, another benefit, it also avoids the bias from the long context with length much larger than $L$ and the case when the question is at the end of the prompt.
We do not compare the LagKV score to the attention weights here. The attention weights vary on different incoming queries. But our scoring method does not depend on the query states or the tokens after the next joint chunk. It mainly finds the tokens that are not coherent to the next chunk and keep them in the cache. As in KIVI Liu et al. (2024c) quantization method, we need a rightful mean to find the correct variance and then prune the small ones. However, we use this strategy to evict tokens instead of quantizing them.
To calculate the compression ratio, we set the token retention ratio as $r$ in each partition. In the partition chunk, only $rL$ tokens will be kept and others are evicted. Therefore, the compression ratio $C$ for the token sequence length $L_{s}\geq S+2L$ can be expressed as:
$L_{R}=S+rL(\lfloor\frac{L_{s}-S}{L}\rfloor-1)+L+Mod(L_{s}-S,L)$ (12)
@@
Where $L_{R}$ is the length of the KV cache after compression. For the case $L_{s}<S+2L$, the compression ratio is zero.
3 Comparisons
3.1 Base Models
@@
We employ two open-source base models: Llama-3.1-8B-Instruct Grattafiori et al. (2024) and Qwen2.5-7B-Instruct Qwen et al. (2025). These models are main stream LLMs with moderate size and both leverage the GQA Ainslie et al. (2023) technique to reduce the KV cache size.
3.2 Results of RULER
We use the RULER Hsieh et al. (2024) benchmark to compare our approach to SnapKV Li et al. (2024) and StreamingLLM Xiao et al. (2023) in various compression ratios. To fairly compare different methods, we integrate our approach into the framework KVPress NVIDIA (2024) and adapt their versions of other approaches. This framework applies compression without question to avoid the query-aware bias. In this task, we set the lag size to be $L=128$ for LagKV and the retention ratio of each recursive window will be adaptively changed by Eq. 13 for different compression ratios.
The results are present in Table 1 and  2 with best scores of each compression ratio shown in bold. The average scores of RULER tasks show that LagKV outperforms SnapKV and StreamingLLM across all compression ratios.
Table 1: RULER-16K Results of Llama-3.1-8B-Instruct
Comp. Ratio Method SK1 SK2 SK3 MK1 MK2 MK3 MV MQ VT CWE FWE QA1 QA2 AVERAGE
0.0 FullKV 100.0 100.0 100.0 97.4 100.0 100.0 100.0 98.2 100.0 90.2 87.5 75.7 54.7 92.6
0.25 SnapKV 100.0 100.0 33.3 98.7 83.3 63.9 97.9 98.2 94.8 85.3 90.2 64.9 46.9 81.3
StreamingLLM 72.5 74.7 72.5 79.2 86.7 66.7 72.7 75.0 90.5 0.1 87.1 75.7 43.8 69.0
LagKV 100.0 100.0 95.7 97.4 96.7 56.9 99.4 98.5 100.0 88.4 89.0 74.3 50.0 88.2
0.5 SnapKV 100.0 94.2 15.9 93.5 48.3 15.3 77.9 87.8 94.8 72.3 85.5 44.6 37.5 66.8
StreamingLLM 47.2 46.0 46.4 53.2 50.0 44.4 48.5 52.4 69.5 1.6 83.1 75.7 35.9 50.3
LagKV 100.0 98.8 88.4 98.7 81.7 13.9 97.9 98.5 98.7 65.7 86.3 66.2 45.3 80.0
0.75 SnapKV 93.4 79.3 4.3 52.0 26.7 1.4 33.5 35.7 83.0 17.1 77.2 28.4 26.6 43.0
StreamingLLM 28.6 21.8 20.3 33.8 26.7 23.6 23.2 27.1 43.6 0.9 80.4 33.8 29.7 30.3
LagKV 100.0 98.8 46.4 90.9 33.3 1.4 86.2 92.4 96.1 10.9 73.3 46.0 42.2 62.9
0.875 SnapKV 85.7 43.7 4.3 26.0 15.0 1.4 17.1 14.3 61.3 1.9 66.3 18.9 26.6 29.4
StreamingLLM 12.1 13.8 11.6 26.0 16.7 15.3 12.3 13.4 20.7 0.9 75.3 29.7 29.7 21.3
LagKV 95.6 77.0 5.8 75.3 8.3 1.4 70.6 80.8 89.8 1.5 62.0 29.7 37.5 48.9
Table 2: RULER-16K Results of Qwen2.5-7B-Instruct
Comp. Ratio Method SK1 SK2 SK3 MK1 MK2 MK3 MV MQ VT CWE FWE QA1 QA2 AVERAGE
0.0 FullKV 100.0 100.0 100.0 99.2 99.1 94.2 94.3 100.0 99.0 79.9 93.2 72.6 48.2 90.8
0.25 SnapKV 88.2 90.6 4.5 44.5 57.1 50.0 39.8 44.9 92.2 80.1 92.8 62.9 42.0 60.7
StreamingLLM 76.3 72.5 75.9 78.9 79.5 67.5 71.1 74.4 76.8 74.1 89.3 69.3 36.6 72.5
LagKV 100.0 99.3 86.6 98.4 88.4 24.2 93.9 99.4 99.0 79.3 92.1 66.1 45.5 82.5
0.5 SnapKV 86.8 64.5 3.6 24.2 28.6 9.2 21.9 23.0 90.8 77.8 92.3 40.3 35.7 46.1
StreamingLLM 50.7 42.8 52.7 52.3 45.5 45.8 48.8 51.2 61.5 72.3 88.3 72.6 33.0 55.2
LagKV 100.0 97.8 48.2 98.4 54.5 3.3 93.9 95.3 98.6 74.4 89.5 58.1 42.9 73.5
0.75 SnapKV 82.2 18.8 3.6 13.3 10.7 4.2 13.2 12.2 79.7 64.4 89.3 27.4 27.7 34.4
StreamingLLM 25.0 20.3 22.3 28.9 25.9 20.0 24.4 25.8 35.8 66.9 82.3 32.3 24.1 33.4
LagKV 99.3 87.7 8.9 85.2 6.2 0.8 86.2 83.5 95.6 43.8 69.5 39.5 29.5 56.6
0.875 SnapKV 69.7 8.0 3.6 14.1 6.2 0.8 11.6 11.4 55.1 47.7 79.2 19.4 22.3 26.9
StreamingLLM 11.2 13.8 14.3 18.8 15.2 12.5 13.4 14.0 20.7 56.1 78.1 21.0 18.8 23.7
LagKV 99.3 62.3 3.6 56.2 0.0 0.8 60.2 50.6 93.0 18.8 55.2 29.0 22.3 42.4
4 Ablations
In this section, we fix the sink size to $S=16$ and vary the lag size $L$ and retention ratio $r$. The values of $L$ will be $L=128$, $512$ and $1024$. The values of $r$ will be 2×2$\times$2 ×, 4×4$\times$4 ×, 6×6$\times$6 × and 8×8$\times$8 × which correspond to $r=0.5$, $0.25$, $0.167$, and $0.125$ respectively. Aslo, we will alter the prefilling method to prove the stability of our approach and scoring method for the validity of lag information.
@@ 4 Ablations
Datasets. We use the facility in Yuan et al. (2024) to extensively test our method. It contains two benchmarks: LongBench Bai et al. (2024) and Needle-in-a-HaystackTest with Passkey-Retrieval in a Paul Graham Essays background Kamradt (2023); Mohtashami and Jaggi (2023). We only test the 64-digit passkey retrieval task which is much more challenging. And because we are using a recursive and evicting compression strategy, it’s easier to illustrate some insights with the partial match score other than the exact one in their report. Therefore, the default needle score will be the partial score throughout the work unless otherwise specified. The main result of this ablation is Table 3.
Prefill stage. By default, like many other compression methods, compression begins after prefill completes for each layer. This is an efficient and accurate approach—preserving both the KV cache values and the first generated token (FGT) while reducing KV cache size. However, since we lack a reliable benchmark for long-context and long-generation scenarios, we will extend the passkey retrieval task by enabling chunk-by-chunk compression during prefill. This will help us evaluate how compression impacts long-generation performance, especially the FGT. Also, this chunked prefilling method will be useful for extreme long context processing.
@@
Llama-3.1-8B-Instruct FullKV 40.71 37.90 28.29 68.49 68.00 58.70 47.44 99.44
@@
Qwen-2.5-7B-Instruct FullKV 41.62 45.00 26.41 68.91 100.00 63.60 51.53 100.00